# Quantum tomography with $\tau$ leptons at the FCC-ee Entanglement, Bell inequality violation, $\sin\theta_{W}$ and anomalous copulings

### v1

@@ 2.1 Quantum tomography at work
in which $\cos\theta^{\pm}_{i}$ are the projections of the $\pi^{\pm}$ momentum direction on the $\{{\bf n},{\bf r},{\bf k}\}$ basis, as computed in the rest frame of the decaying $\tau^{\pm}$. Figures 2.1 and 2.2 show the normalized distributions obtained for the values of the $\operatorname{B}_{i}$ and $\operatorname{C}_{ij}$ from the actual Monte Carlo simulation we have used in this work. The central values of the Gaussian fits—dashed lines—indicate the values of the coefficients themselves.
Figure 2.1: Distributions obtained for the coefficients $\operatorname{B}_{i}^{\pm}$ in the Monte Carlo simulation of the $\tau$-lepton pairs decaying into a charged pion plus a neutrino, the momentum of which is reconstructed. The dotted curves represent the Gaussian fit of the histograms.
@@
Figure 2.2: Distribution of the values of the coefficients $\operatorname{C}_{ij}$ in the Monte Carlo simulation of the $\tau$-lepton pairs decaying into a charged pion plus a neutrino, the momentum of which is reconstructed. The dotted curves represent the Gaussian fit of the histograms.
@@ 2.2 Entanglement and Bell inequality violation
The Horodecki’s condition in Eq. (2.8) may show a bias toward positive values when evaluated numerically on samples with a restricted statistics. The issue was addressed in [10] by correcting for it in the analysis of the violation of Bell inequalities in top-quark production at the LHC, a case in which the number of events is limited and $\mathfrak{m}_{12}$ is biased. In this work, as in [11], such a correction is not necessary because, as shown in Fig 2.3, the bias is negligible for samples containing at least $10^{4}$ events.
@@ 2.3 Monte Carlo simulation
We generate 10 million events, containing each a $\tau$ lepton pair decaying into two opposite charged pions plus the $\tau$ neutrinos, by means of MadGraph5 [12] and the TauDecay library [13]. The events are generated at the tree level in the electroweak interaction and, therefore, the cross section and the remaining observables are expected to assume their tree-level values.
To make the simulation closer to the analysis with actual data, we replace the Monte Carlo truth $\tau$ lepton momenta with those obtained from the neutrino momenta reconstruction. We then sort the events into 50 independent samples (each corresponding to a pseudo-experiments with $2\times 10^{5}$ events) and reconstruct the $\operatorname{B}^{\pm}_{i}$ and $\operatorname{C}_{ij}$ coefficients for each sample. The resulting Gaussian distributions, shown in Figures 2.1 and 2.2, give the central values and the statistical uncertainties of these parameters.
To mimic real data even further, we repeat the analysis by modelling the efficiency and uncertainty in the charged pion tracks and vertex determination of the detector prior to performing the neutrino momenta reconstruction. More details are given below.
@@ 2.3.2 Detector response and initial state radiation
We model the detector resolution with the following uncertainties:
$\frac{\sigma_{p_{T}}}{p_{T}}=3\times 10^{-5}\oplus 0.6\times 10^{-3}\frac{p_{T}}{\rm GeV}\quad\text{and}\quad\sigma_{\theta,\phi}=0.1\times 10^{-3}\,\text{rad}$ (2.15)
@@
$\sigma_{b}=3\,\mu{\rm m}\oplus\frac{15\,\mu{\rm m}}{\sin^{2/3}\Theta}\frac{\rm GeV}{p_{T}}$ (2.16)
for the impact parameters and for the reconstruction of the vector of closest approach. These values are taken from the envisaged IDEA tracking detector [1]; a constant smearing originates in imperfections in the detector or anomalies in signal collection, whereas the term scaling with the momentum is a noise that comes from the imperfection of the readouts. In the simulation we retain the leading contributions in the quoted uncertainties and account for the related detector effects by performing a Gaussian smearing of the Monte Carlo truth pion momenta and closest approach vector taking the nominal resolutions above as standard deviations.
@@
Figure 2.4: Distribution of the $\tau$-lepton pair events in invariant mass after including the initial state radiation of the electrons. The last four bins, from 89 to 91 GeV, contain 99% of the events.
@@ 3.1 Theoretical quantum tomography: analytic results
Let us first look at the analytic results for the parton level process $e^{+}e^{-}\to Z\to\tau^{+}\tau^{-}$. This study already provides most of the information about the actual physical process, that is, the one inclusive of the hadronic decays of the $\tau$ leptons.
@@
Figure 3.2: Values of the concurrence $\mathscr{C}$ and the Horodecki’s condition $\mathfrak{m}_{12}$ as functions of the scattering angle in the parton level process $e^{+}e^{-}\to\tau^{+}\tau^{-}$. The presence of entanglement is signaled by $\mathcal{C}>0$ and the violation of Bell inequalities by $\mathfrak{m}_{12}>1$.
The values of the concurrence $\mathscr{C}$ and the Horodecki’s condition $\mathfrak{m}_{12}$ for energies that range from the production threshold up to 350 GeV are shown in Fig. 3.1. The result for $\sqrt{s}=10$ GeV have been already discussed in [11] and applied to superKEKB and the physics at Belle II. Here we study the region at the $Z$-boson resonance. In Fig. 3.2 the C M energy is set at $\sqrt{s}=91.19$ GeV and the concurrence and the Horodecki’s condition are shown as functions of the scattering angle $\Theta$, formed by the $e^{+}$ and $\tau^{+}$ directions.
As for the polarization of the $\tau$ pair, we find $\operatorname{B}_{i}^{+}=\operatorname{B}_{i}^{-}=\operatorname{B}_{i}$ for $i=r,\,n,\,k$. Contrary to most of the other process that have been utilized in studies of entanglement, we have non vanishing $\operatorname{B}_{i}$ coefficients at the leading order even though the initial beam is made of unpolarized electrons and positrons. This feature is due to the parity violating electroweak interactions and makes the study of the $\tau$ leptons at this energy particularly interesting. The two plots in Fig. 3.3 show the behavior of the coefficients $\operatorname{B}_{i}^{\pm}$ as functions of the scattering angle $\Theta$ at $\sqrt{s}=91.19$ GeV, as well as in a broader energy range.
@@ 3.2 Tomography with Monte Carlo simulation
As explained in Section 2, we generate a total of $10^{7}$ events containing $\tau$ pairs that result in two opposite charged pions and missing energy in the final states. For each of the events we reconstruct the rest frames of the $\tau^{\pm}$ leptons and compute the projection of their spin, indicated by the directions of the pions, on the $\{{\bf r},{\bf n},{\bf k}\}$ basis in these frames of reference. The statistical uncertainty is estimated by running 50 independent pseudo-experiments each consisting of a sample of 200 000 events drawn from the full dataset of $10^{7}$ Monte Carlo generated events.
Omitting, for the moment, the effects of ISR and finite detector resolution, the procedure leads to the following determination of the coefficients $\operatorname{C}_{ij}$ and their statistical uncertainty:
@@ 3.2.1 Monte Carlo simulation with detector and ISR effects
$\boxed{\mathfrak{m}_{12}=1.239\pm 0.017\,.}$ (3.12)
The significance of the Bell inequality violation is about 14 standard deviations. Equations Eqs. (3.11)–(3.12) are the main result of the present work. The significance can be increased by means of kinematic cuts that select events with scattering angles close to $\Theta=\pi/2$, for which the values of $\mathcal{C}$ and $\mathfrak{m}_{12}$ are larger (see Fig. 3.2).
@@ 3.2.2 Background and systematic uncertainties
The absolute value of the difference between the central values of the concurrence obtained in the Monte Carlo simulation without, see Eq. (3.7), and with detector effects and ISR, see Eq. (3.11), is
$\delta\mathscr{C}|_{\rm syst}=\left|0.481-0.485\right|=0.004\,;$ (3.13)
@@
These differences provide an estimate of the upper bound in the systematic errors of these two observables. They are subdominant with respect to the statistical errors in Eq. (3.11) and Eq. (3.12).
To find the expected uncertainty at the FCC-ee—where about $10^{9}$ events are to be collected in four years of operation—we must rescale our uncertainties obtained by running pseudo-experiments containing 200 000 event each by a factor $10\,\sqrt{50}$. This rescaling will make the statistical uncertainties smaller than the upper limits on the systematic errors, which at that point will come to dominate the overall uncertainty unless reduced and brought under control. We expect that indeed this will be the case and that the systematic error will be under control after the operating of the machine and the detectors are comprehensively understood.
@@ 3.3 Polarizations
From the Monte Carlo simulation with detector effects and ISR included, we take the polarization to be the arithmetic average and obtain
$\boxed{\langle P\rangle_{\tau}=\frac{1}{2}(\operatorname{B}_{k}^{+}+\operatorname{B}_{k}^{-})=0.2203\pm 0.0044\,,}$ (3.15)
while $\operatorname{B}_{n}^{\pm}$ and $\operatorname{B}_{r}^{\pm}$ are consistent with their vanishing analytic value, and the overall sign depends on the frame of reference. The polarization in Eq. (3.15) corresponds to the tree-level value, the measured value being $0.1476\pm 0.0108$ [4].
@@ 3.3.2 Weinberg angle
$\boxed{\sin^{2}\theta_{W}=0.2223\pm 0.0006\,,}$ (3.19)
@@
Assuming that the systematic uncertainties will eventually be under control, we rescale the uncertainty of our sample to account for the expected FCC-ee yield and obtain an uncertainty of $7.8\times 10^{-6}$ on the determination of the tree-level Weinberg angle. This uncertainty is two order of magnitude better than the current value [19]. The report [1] cites uncertainties in the Weinberg angle determination between $10^{-5}$ and $10^{-6}$ as one of the goals for the FCC-ee machine.
@@ 3.4.1 Observables
For the present case, we use the entries of the polarization density matrix to define three observables that provide the means to best constrain the parameters in Eq. (3.20): one observable is the concurrence $\mathscr{C}$ defined in Eq. (2.7) and it measures the entanglement in the spin states of the produced $\tau$ pairs, another— specific to the CP-violating contributes—is related to triple products involving one momentum and the spin vectors of the $\tau$ leptons:
@@ 3.4.2 Limits
Figure 3.5: $\chi^{2}$ test for the form factor $F_{1}^{V}(m_{Z}^{2})$ (left) and $F_{1}^{A}(m_{Z}^{2})$ (right). Both limits are obtained by means of the cross section.
For a sample of 200.000 events, the expected value and uncertainty on the cross section is $(22.83\pm 0.05)$ pb. For the same sample, from Section 3.2.1, we find an uncertainty on $\operatorname{C}_{ij}$, and therefore on $\mathscr{C}_{odd}$, of at most 0.009, and on the concurrence $\mathscr{C}$ of 0.006. We use these uncertainties in the determination of the limits on the form factors with the understanding that they can be improved by rescaling the number of events to that actually expected at the FCC-ee.
@@
The bounds on each coupling can be extracted from Fig. 3.4 and 3.5 under the assumption that the parameters are independent. With that, we neglect possible correlations. The values obtained from the 95% confidence interval are reported in Table 3.1.
           $\mathscr{O}_{a}$            $\sigma_{a}$            form factor limits
$\mathscr{C}$ 0.006 $-0.002\leq F_{2}(m_{Z}^{2})\leq 0.003$
$\mathscr{C}_{odd}$ 0.009 $-0.001\leq F_{3}(m_{Z}^{2})\leq 0.001$
$\sigma_{T}$ 0.05 pb $-0.009\leq C_{1}^{V}\leq 0.010$
$\sigma_{T}$ 0.05 pb $-0.001\leq C_{1}^{A}\leq 0.001$
Table 3.1: Bounds obtained at the 95% confidence level for the form factors, as shown in Figures 3.4 and 3.5, neglecting correlations. The values refer to a sample of 200 000 events; limits for larger samples can be obtained by rescaling these values by the square root of the ratio of the corresponding numbers of events.
@@
The limit on $F_{3}(m_{Z}^{2})$ in Table 3.1 corresponds to a limit on the CP violating weak dipole moment of $5.6\times 10^{-18}$ e cm for the $\tau$ lepton. The corresponding limit from LEP is equal to $1.4\times 10^{-17}$ [26, 27]. These limits are comparable since they are obtained with a similar number of events and utilize similar observables built out of a CP-odd triple product. Rescaling our limits to the FCC-ee luminosity, we find $|F_{3}(m_{Z}^{2})|\leq 2.0\times 10^{-19}$, which is one order of magnitude better than the current limit: $3.6\times 10^{-18}$ e cm [19].
Table 3.1 shows the bounds obtained with pseudo-experiments containing each 200 000 events. Corresponding projections base on the full dataset of $10^{9}$ expected at the FCC-ee can be derived (considering only the statistical errors) upon a rescaling by a factor $10\sqrt{50}$.

### v2

# Quantum tomography with $\tau$ leptons at the FCC-ee Entanglement, Bell inequality violation, $\sin\theta_{W}$ and anomalous couplings

@@ 2.1 Quantum tomography at work
in which $\cos\theta^{\pm}_{i}$ are the projections of the $\pi^{\pm}$ momentum direction on the $\{{\bf n},{\bf r},{\bf k}\}$ basis, as computed in the rest frame of the decaying $\tau^{\pm}$. Figures 2.1 and 2.2 show the relative frequencies of the values obtained for the involved cosines and cosine products from the actual Monte Carlo simulation we have used in this work. All the histograms are normalized to the total number, $10^{7}$, of simulated events. The average values of these these quantities, proportional to the $\operatorname{B}_{i}$ and $\operatorname{C}_{ij}$ coefficients, encode the tomographic information and do not vanish provided the corresponding histograms are not symmetric with respect to the zero value.
Figure 2.1: Relative frequencies of the values obtained for $\cos\theta_{i}^{\pm}$, $i=n,r,k$, in the Monte Carlo simulation of the $\tau$-lepton pairs decaying into a charged pion plus a neutrino, the momentum of which is reconstructed. The histograms are normalized to the total number of events.
@@
Figure 2.2: Relative frequencies of the values of the products $\cos\theta^{+}_{i}\cos\theta^{-}_{j}$, $i,j=n,r,k$, obtained with the Monte Carlo simulation of the $\tau$-lepton pairs decaying into a charged pion plus a neutrino, the momentum of which is reconstructed. The histograms are normalized to the total number of events.
@@ 2.2 Entanglement and Bell inequality violation
The Horodecki’s condition in Eq. (2.8) may show a bias toward positive values when the eigenvalues are evaluated numerically on samples with a restricted statistics. The issue was addressed in [10] by correcting for it in the analysis of the violation of Bell inequalities in top-quark production at the LHC, a case in which the number of events is limited and $\mathfrak{m}_{12}$ is biased. In this work, as in [11], such a correction is not necessary because, as shown in Fig 2.3, the bias is negligible for samples containing at least $10^{4}$ events.
@@ 2.3 Monte Carlo simulation
We generate 10 million events, containing each a $\tau$ lepton pair decaying into two opposite charged pions plus the $\tau$ neutrinos, by means of MadGraph5 [12] and the TauDecay library [13]. No cuts other than those in the default run card were applied. The events are generated at the tree level in the electroweak interaction and, therefore, the cross section and the remaining observables are expected to assume their tree-level values. The generated number of events provides an adequate benchmark to probe the capabilities of the FCC-ee to perform quantum information analyses. The expected uncertainty at the FCC-ee—where about $10^{9}$ events are to be collected in four years of operation with an integrated luminosity of 150 ab-1—can be obtained upon a rescaling by a factor $10\,\sqrt{50}$ of the statistical uncertainties indicated by our benchmark.
To make the simulation closer to the analysis with actual data, we replace the Monte Carlo truth $\tau$ lepton momenta with those obtained from the neutrino momenta reconstruction. We then sort the events into 50 independent samples (each corresponding to a pseudo-experiments with $2\times 10^{5}$ events equivalent to an effective integrated luminosity of 17.6 fb-1) and reconstruct the $\operatorname{B}^{\pm}_{i}$ and $\operatorname{C}_{ij}$ coefficients for each sample.
To mimic real data even further, we repeat the analysis by modeling the efficiency and uncertainty in the charged pion tracks and vertex determination of the detector prior to performing the neutrino momenta reconstruction. More details are given below.
@@ 2.3.2 Detector response and initial state radiation
We model the detector resolution with two sets of uncertainties meant to mimic different systematic errors in the reconstructions. The single smearing of momenta and tracks simulate the detector response and limitations while the difference in the results obtained by means of the two sets is a measure of the experimental systematic uncertainty. We have chosen the following two sets
$\frac{\sigma_{p_{T}}}{p_{T}}=3\times 10^{-5}\oplus 0.3\times 10^{-3}\frac{p_{T}}{\rm GeV}\quad\text{and}\quad\sigma_{\theta,\phi}=0.1\times 10^{-3}\,\text{rad}$ (2.15)
or
$\frac{\sigma_{p_{T}}^{\prime}}{p_{T}}=3\times 10^{-5}\oplus 0.6\times 10^{-3}\frac{p_{T}}{\rm GeV}\quad\text{and}\quad\sigma_{\theta,\phi}=0.1\times 10^{-3}\,\text{rad}$ (2.16)
@@
$\sigma_{b}=3\,\mu{\rm m}\oplus\frac{15\,\mu{\rm m}}{\sin^{2/3}\Theta}\frac{\rm GeV}{p_{T}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\quad\text{or}\quad\sigma_{b}^{\prime}=5\,\mu{\rm m}\oplus\frac{15\,\mu{\rm m}}{\sin^{2/3}\Theta}\frac{\rm GeV}{p_{T}}}$ (2.17)
for the impact parameters and for the reconstruction of the vector of closest approach. These sets of values are taken from the envisaged IDEA tracking detector [1, 16]. The specific choice we made provides a useful benchmark to guide our analysys. A constant smearing originates in imperfections in the detector or anomalies in signal collection, whereas the term scaling with the momentum is a noise that comes from the imperfection of the readouts. In the simulation we retain the leading contributions in the quoted uncertainties and account for the related detector effects by performing a Gaussian smearing of the Monte Carlo truth pion momenta and closest approach vector taking the nominal resolutions above as standard deviations.
@@
Figure 2.4: Distribution of the $\tau$-lepton pair events in invariant mass after including the initial state radiation of the electrons. The last four bins, from 89 to 91 GeV, contain 99% of the events. The count in each bin is normalized to the total number of events.
@@ 3.1 Theoretical quantum tomography: analytic results
Let us first look at the analytic results for the parton level process $e^{+}e^{-}\to\tau^{+}\tau^{-}$. This study already provides most of the information about the actual physical process, that is, the one inclusive of the hadronic decays of the $\tau$ leptons.
@@
Figure 3.2: Values of the concurrence $\mathscr{C}$ and the Horodecki’s condition $\mathfrak{m}_{12}$ as functions of the scattering angle in the parton level process $e^{+}e^{-}\to\tau^{+}\tau^{-}$ for $\sqrt{s}=91.19$ GeV. The presence of entanglement is signaled by $\mathcal{C}>0$ and the violation of Bell inequalities by $\mathfrak{m}_{12}>1$.
The values of the concurrence $\mathscr{C}$ and the Horodecki’s condition $\mathfrak{m}_{12}$ for energies that range from the production threshold up to 350 GeV are shown in Fig. 3.1. The plot shows the presence of three distinct regimes as we vary the CM energy. At low energy the process is dominated by the photon exchange, yielding the maximal value of the concurrence and of the Bell inequality violation for $\Theta=\pi/2$. At energies close to the $Z$ boson resonance the process is instead dominated by this particle. In Fig. 3.1 this region corresponds to the thin slice centered around $\sqrt{s}\simeq 90$ GeV. At higher energies both the particles equally contribute to the process and the maximal concurrence and Bell inequality violation are achieved for a different value of the scattering angle due to the different interplay among the couplings. In between different regions, interference effects reduce the entanglement.
The result for $\sqrt{s}=10$ GeV have been already discussed in [11] and applied to superKEKB and the physics at Belle II. Here we study the region at the $Z$-boson resonance.
In Fig. 3.2 the CM energy is set at $\sqrt{s}=91.19$ GeV and the concurrence and the Horodecki’s condition are shown as functions of the scattering angle $\Theta$, formed by the $e^{+}$ and $\tau^{+}$ directions.
As for the polarization of the $\tau$ pair, we find $\operatorname{B}_{i}^{+}=\operatorname{B}_{i}^{-}=\operatorname{B}_{i}$ for $i=r,\,n,\,k$. Contrary to most of the other process that have been utilized in studies of entanglement, we have non vanishing $\operatorname{B}_{i}$ coefficients at the leading order even though the initial beam is made of unpolarized electrons and positrons. This feature is due to the parity violating electroweak interactions and makes the study of the $\tau$ leptons at this energy particularly interesting. The two plots in Fig. 3.3 show the behavior of the coefficients $\operatorname{B}_{i}^{\pm}$ as functions of the scattering angle $\Theta$ at $\sqrt{s}=91.19$ GeV, as well as in a broader energy range. The characteristic pattern in the CM energy dependence is due to the interference between the photon and $Z$ boson contributions. The angular dependence is brought about by the interplay between the axial and vector couplings and the scattering angle.
@@ 3.2 Tomography with Monte Carlo simulation
As explained in Section 2, we generate a total of $10^{7}$ events containing $\tau$ pairs that result in two opposite charged pions and missing energy in the final states. For each of the events we reconstruct the rest frames of the $\tau^{\pm}$ leptons and compute the projection of their spin, indicated by the directions of the pions, on the $\{{\bf r},{\bf n},{\bf k}\}$ basis in these frames of reference.
The distributions in Figs. 2.1 and 2.2 of the coefficients obtained with the full Monte Carlo sample are not necessarily Gaussian and the skewness is what actually determines a non-vanishing average value. In order to infer a standard error and the best estimate for the value of each coefficient we then use the central limit theorem, which ensures that the sample means are normally distributed around the mean of the full Monte Carlo sample, regardless of the actual distribution of the full population. By computing the mean of the sample means and the related variance for each parameter, we are then ensured that the obtained values are faithful estimates of the corresponding true values and standard error. Having generated $10^{7}$ events, we opted to use 50 samples of 200 000 events each to implement the procedure (corresponding to an integrated luminosity of 17.6 fb-1). We checked that the population mean of each parameter falls within the $1\sigma$ confidence interval indicated by our method.
Omitting, for the moment, the effects of ISR and finite detector resolution, the outlined procedure leads to the following determination of the coefficients $\operatorname{C}_{ij}$ and their statistical uncertainty:
@@ 3.2.1 Monte Carlo simulation with detector and ISR effects
$\mathfrak{m}_{12}=1.239\pm 0.017\,.$ (3.12)
@@ 3.2.2 Background and systematic uncertainties
Theoretical systematic errors are either very small (those coming from NLO contributions or the uncertainty in the value of $m_{Z}$) or not present (those originating from different parton distribution functions which are not utilized in the first place) because of the EW nature of the process under consideration. A realistic determination of the experimental systematic uncertainties is not possible until the machine and the detectors are comprehensively understood. Yet we can partially estimate the size of the systematic errors affecting the concurrence and the Horodecki condition by accounting for ISR and the projected detector effects. To this end, we first compare the absolute value of the difference between the central values of the concurrence obtained in the Monte Carlo simulation without, see Eq. (3.7), and with detector effects and ISR, see Eq. (3.11)
$\delta\mathscr{C}|_{\rm syst}=\left|0.4815-0.4805\right|=0.0010\,,$ (3.13)
@@
These differences account for the systematic uncertainty in the beam energy and are subdominant with respect to the statistical errors in Eq. (3.11) and Eq. (3.12).
An additional contribution to the systematic errors can be obtained by using the two sets of uncertainties in the smearing procedure introduced in Section 2.3.2. These yield
$\delta\mathscr{C}|_{\rm syst}=\left|0.4807-0.4805\right|=0.0002\,,$ (3.15)
and
$\delta\mathfrak{m}_{12}|_{\rm syst}=\left|1.232-1.230\right|=0.002\,.$ (3.16)
These contributions are, again, subdominant with respect to the statistical errors in Eq. (3.11) and Eq. (3.12). They become important as we rescale our statistical uncertainties toward the FCC target luminosity and will eventually come to dominate.
Including now also the systematic errors estimated above we find that
$\boxed{\mathscr{C}=0.4805\pm 0.0063|_{\rm stat}\pm 0.0012|_{\rm syst}\,,}$ (3.17)
and the Horodecki’s condition gives
$\boxed{\mathfrak{m}_{12}=1.239\pm 0.017|_{\rm stat}\pm 0.008|_{\rm syst}\,,}$ (3.18)
where the systematic errors where added linearly.
Equations Eqs. (3.17)–(3.18) are the main result of the present work. The overall significance of the Bell inequality violation (obtained with a benchmark luminosity of 17.6 fb-1) is about 13 standard deviations once the errors are added in quadrature. The expected significance with 150 ab-1 of data can be estimated by retaining only the systematic error in Eq. (3.18) and it is about 30 standard deviations.
These significances can be increased by means of kinematic cuts that select events with scattering angles close to $\Theta=\pi/2$, for which the values of $\mathcal{C}$ and $\mathfrak{m}_{12}$ are larger (see Fig. 3.2).
@@ 3.3 Polarizations
From the Monte Carlo simulation with detector effects and ISR included (corresponding to data collected with a benchmark integrated luminosity of 17.6 fb-1), we take the polarization to be the arithmetic average and obtain
$\boxed{\langle P\rangle_{\tau}=\frac{1}{2}(\operatorname{B}_{k}^{+}+\operatorname{B}_{k}^{-})=0.2203{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm 0.0044|_{\rm stat}\pm 0.0008|_{\rm syst}}\,,}$ (3.19)
while $\operatorname{B}_{n}^{\pm}$ and $\operatorname{B}_{r}^{\pm}$ are consistent with their vanishing analytic value, and the overall sign depends on the frame of reference. The systematic error is obtained as discussed in the previous section. The polarization in Eq. (3.19) corresponds to the tree-level value, consistently with the input value used for the Weinberg angle—see Sec. 3.3.2—whereas the current measured value is $0.1476\pm 0.0108$ [4].
@@ 3.3.2 Weinberg angle
$\boxed{\sin^{2}\theta_{W}=0.2223{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm 0.0006|_{\rm stat}\pm 0.0001|_{\rm syst}}\,,}$ (3.23)
@@
Rescaling the statistical uncertainty of our sample (obtained with a benchmark luminosity of 17.6 fb-1) to account for the expected FCC-ee yield, we obtain an uncertainty of $1\times 10^{-4}$, completely dominated by the systematics. This uncertainty is of the same order of magnitude of the current value [20]. The report [1] cites uncertainties in determination of the Weinberg angle between $10^{-5}$ and $10^{-6}$ as one of the goals for the FCC-ee machine—such a value can only be achieved if the systematic error indicated by our estimate is reduced.
@@ 3.4.1 Observables
For the present case, we use the entries of the polarization density matrix to define three observables used to constrain the parameters in Eq. (3.24): one observable is the concurrence $\mathscr{C}$ defined in Eq. (2.7) and it measures the entanglement in the spin states of the produced $\tau$ pairs, another— specific to the CP-violating contributes—is related to triple products involving one momentum and the spin vectors of the $\tau$ leptons:
@@ 3.4.2 Limits
Figure 3.5: $\chi^{2}$ test for the form factor $C_{1}^{V}$ (left) and $C_{1}^{A}$ (right). Both limits are obtained by means of the cross section.
For a sample of 200.000 events, the expected value and uncertainty on the cross section is $(22.83\pm 0.05)$ pb. For the same sample, from Section 3.2.1, we find an uncertainty on $\operatorname{C}_{ij}$, and therefore on $\mathscr{C}_{odd}$, of at most 0.009, and on the concurrence $\mathscr{C}$ of 0.006. We use these uncertainties in the determination of the limits on the form factors. For the number of events actually expected at the FCC-ee, we rescale the statistical errors by $10\sqrt{50}$, which therefore becomes negligible, and use the systematic errors as the whole uncertainties, namely, a 1 per mille uncertainty in the cross section, $0.006$ for the operator $\mathscr{C}_{odd}$ and $0.001$ for the concurrence $\mathscr{C}$.
@@
The bounds on each coupling can be extracted from Fig. 3.4 and 3.5 in the case of 200 000 events and from similar figures in the case of $10^{9}$ events. The corresponding values obtained from the 95% confidence interval are reported in Table 3.1. Though the quoted bounds neglect possible correlations among the parameters, they are sufficient to support our claim that quantum information observables can be used to effectively constrain new physics effects.
   $\mathscr{O}_{a}$       $\sigma_{a}^{I}$    limits I (L = 17.6 fb-1)       $\sigma_{a}^{II}$    limits II (L = 150 ab-1)
$\mathscr{C}$ 0.006 $-0.002\leq F_{2}(m_{Z}^{2})\leq 0.003$ 0.001 $-0.001\leq F_{2}(m_{Z}^{2})\leq 0.001$
$\mathscr{C}_{odd}$ 0.009 $-0.001\leq F_{3}(m_{Z}^{2})\leq 0.001$ 0.006 $-0.0004\leq F_{3}(m_{Z}^{2})\leq 0.0005$
$\sigma_{T}$ 0.05 pb $-0.009\leq C_{1}^{V}\leq 0.010$ 0.02 pb $-0.004\leq C_{1}^{V}\leq 0.004$
$\sigma_{T}$ 0.05 pb $-0.001\leq C_{1}^{A}\leq 0.001$ 0.02 pb $-0.0004\leq C_{1}^{A}\leq 0.0004$
Table 3.1: Bounds obtained at the 95% confidence level for the form factors, as shown in Figures 3.4 and 3.5, neglecting correlations. The values I refer to a sample of 200 000 events, corresponding to a luminosity $L$ of 17.6 fb-1; the values II are for the sample of $10^{9}$ events expected at the FCC after 4 years of operation with an integrated luminosity of 150 ab-1. The uncertainty in the latter case is given by the estimated systematic error (we set that for the cross section at 1 per mille.)
@@
The limit on $F_{3}(m_{Z}^{2})$ in Table 3.1 corresponds to a limit on the CP violating weak dipole moment of $5.6\times 10^{-18}$ e cm for the $\tau$ lepton. The corresponding limit from LEP is equal to $1.4\times 10^{-17}$ [27, 28]. These limits are comparable since they are obtained with a similar number of events and utilize similar observables built out of a CP-odd triple product. Rescaling our limits to the FCC-ee luminosity, we find $|F_{3}(m_{Z}^{2})|\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1.1\times 10^{-19}}$ e cm, which is one order of magnitude better than the current limit: $3.6\times 10^{-18}$ e cm [20].
Table 3.1 shows the bounds obtained with pseudo-experiments containing each 200 000 events (second column) and the corresponding projections base on the full dataset of $10^{9}$ expected at the FCC-ee (fourth column), in which only the systematic uncertainty is considered since the statistical errors are negligible upon rescaling by a factor $10\sqrt{50}$.